[Brahms-l] Shift report 20050612 16:00-24:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Sun Jun 12 2005 - 00:01:06 EDT
Shifter leader was Theresa Wilkerson
Shifters were 


1. Machine status ok, except for last hour: STAR magnets tripped at 2310.

2. Run #15411-15424. Useful data taken for #15413-15424. Triggers from Display Scaler program used for #15414-15424. Aprx. 6 beam hours.

3. TOFW Top + Bot: malfunctioning slats, see shift report with accompanying diagram. Also note that a polarization measurement was taken during run #15417.

Special note: MCR has requested that we use lower field settings for our magnets, owing to greater beam intensity observed on Thursday morning while BRAHMS was analyzing FFS data. See shift log for further explanation and specific settings.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sun Jun 12 00:01:42 2005

This archive was generated by hypermail 2.1.8 : Sun Jun 12 2005 - 00:01:54 EDT